# Invariant subspaces and PCA in nearly matrix multiplication time

### v3

@@ A.2 Floating point model
We assume a standard floating point model of computation and borrow its axioms from [Hig02, Chapter 2]. There is a fixed number of bits to represent floating point numbers, specifically, one bit is reserved for the sign $s$, $p$ bits are used for the exponent $e$, and $t$ bits are used for the significand $m$.
@@
From [Hig02, Theorem 2.2], for all real numbers $\alpha$ in the normalized range it holds that
@@
Divisions and multiplications with $1$ and $2$ do not introduce errors (for the latter we simply increase/decrease the exponent). We assume that we also have an implementation of $\sqrt{\cdot}$ such that $\boldsymbol{\mathsf{fl}}(\sqrt{\alpha})=(1+\theta)\sqrt{\alpha}$ where $|\theta|\leq\textbf{{u}}$. From [Hig02, Lemma 3.1], we can bound products of errors as
@@
The above can be extended also for complex arithmetic (see [Hig02, Lemma 3.5]), where the bound becomes $|\theta|\leq O(\textbf{{u}})$, but we will ignore the constant prefactor for simplicity.
@@ A.3 Spectral decomposition, pseudospectrum, and eigenvalue bounds
In the sections that follow we need to bound the (forward) errors on the eigenvalues of matrices under perturbations that are introduced due to the finite precision arithmetic. Such bounds can be derived by the classic Bauer-Fike theorem the following (cf. [BF60, EI98] for more details).
There are many bounds in the literature describing the effect of perturbations on the eigenvalues. We summarize some classic results in the following proposition and refer to Bhatia’s monograph for a detailed overview [Bha07].
@@
Hermitian Hermitian $|\lambda_{i}(\boldsymbol{\mathrm{A}})-\lambda_{i}(\boldsymbol{\mathrm{B}})|\leq\|\boldsymbol{\mathrm{A}}-\boldsymbol{\mathrm{B}}\|$ [Wey12]
Hermitian Non-Hermitian $|\lambda_{i}(\boldsymbol{\mathrm{A}})-\lambda_{i}(\boldsymbol{\mathrm{B}})|\leq O(\log(n))\|\boldsymbol{\mathrm{A}}-\boldsymbol{\mathrm{B}}\|$ [Kah75, Pok81]
@@
Recall some useful properties from the seminal book of Trefethen and Embree [TE05].
Proposition A.1 (Collective results from [TE05]).
@@
    $\Lambda_{\epsilon-\|\boldsymbol{\mathrm{E}}\|}(\boldsymbol{\mathrm{M}})\subseteq\Lambda_{\epsilon}(\boldsymbol{\mathrm{M}}+\boldsymbol{\mathrm{E}})\subseteq\Lambda_{\epsilon+\|\boldsymbol{\mathrm{E}}\|}(\boldsymbol{\mathrm{M}})$ ([TE05, Thm 52.4]),
@@
    Any bounded connected component of $\Lambda_{\epsilon}(\boldsymbol{\mathrm{M}})$ has a nonempty intersection with $\Lambda(\boldsymbol{\mathrm{M}})$, i.e., it contains at least one eigenvalue ([TE05, Thm 4.3]),
@@
    $\bigcup_{i}D(\lambda_{i},\epsilon)\subseteq\Lambda_{\epsilon}(\boldsymbol{\mathrm{M}})\subseteq\bigcup_{i}D(\lambda_{i},\epsilon\kappa_{\boldsymbol{\mathrm{V}}}(\boldsymbol{\mathrm{M}}))$ ([TE05, Thms 4.3 and 52.2]).
@@ A.5 Imported floating point algorithms for fast linear algebra
Before we dive into the details, we recall the concept of “fast matrix multiplication”, pioneered by Strassen [Str69], who showed that two square matrices can be multiplied in $O(n^{\omega})$, where $\omega=\log_{2}7\approx 2.807<3$, in real arithmetic. Since then, the matrix multiplication exponent $\omega$ has been significantly reduced, the record to date being $\omega\leq 2.371552$ [DWZ23, WXXZ24]. In two seminal works [DDHK07, DDH07], Demmel, Dumitriu, Holtz, and Kleinberg, proved that any fast matrix multiplication algorithm can be executed numerically stably in a floating point machine with almost the same arithmetic complexity as in real arithmetic. They also showed that other problems in Numerical Linear Algebra can be reduced to such matrix multiplications, including inversion, LU and QR factorizations, and solving linear systems of equations. Our algorithms build on the existing results for stable fast matrix multiplication [DDHK07], inversion [DDH07], as well as backward-approximate diagonalization and the matrix sign function [BGVKS22]. We import the corresponding results from the aforementioned works in the following theorems.
Theorem A.2 ($\mathsf{MM}$, stable fast matrix multiplication [DDHK07]).
@@
Theorem A.3 ($\mathsf{INV}$, logarithmically-stable fast inversion [DDH07]).
@@
Definition A.3 (Circles of Apollonius, imported from Section 4.1 of [BGVKS22]).
@@
Theorem A.4 ($\mathsf{SGN}$, imported Theorem 4.9 from [BGVKS22]).
@@
Theorem A.5 ($\mathsf{DEFLATE}$, imported Theorem 5.3 of [BGVKS22]).
@@
The algorithm succeeds with probability at least $1-\tfrac{(20n)^{3}\sqrt{\beta}}{\epsilon^{2}}$. The number of arithmetic operations is at most $O(T_{\mathsf{MM}}(n))$, and it requires $O(\log(n)+\log(\tfrac{1}{\epsilon}))$ bits of precision. Internally, the algorithm generates $O(n^{2})$ random numbers to form a $n\times n$ complex Ginibre matrix. [^1] We highlight that our Gaussian sampling oracle in Definition D.1 is slightly different than the one described in [BGVKS22], but it should be straightforward to adapt the analysis using either of the two definitions.
@@
Lastly, we recall the following two results. The first one is for the QR factorization from [DDH07], which we adapt suitably for our analysis, specifically, to compute an orthonormal basis for the column space of a rectangular matrix.
Theorem A.6 (Basis computation, follows from Section 4 in [DDH07]).
@@
We now use the corresponding algorithm of [DDH07] on $\boldsymbol{\mathrm{A}}^{\prime}$. It returns three matrices: an upper triangular matrix $\widetilde{\boldsymbol{\mathrm{R}}}\in\mathbb{R}^{n\times n}$, a matrix $\boldsymbol{\mathrm{W}}\in\mathbb{R}^{m\times n}$ with $\|\boldsymbol{\mathrm{W}}\|\in O(n)$, and a matrix $\boldsymbol{\mathrm{Y}}\in\mathbb{R}^{n\times m}$ with $\|\boldsymbol{\mathrm{Y}}\|\in O(n)$. The matrix $\boldsymbol{\mathrm{\Psi}}=\boldsymbol{\mathrm{I}}-\boldsymbol{\mathrm{W}}\boldsymbol{\mathrm{Y}}^{\top}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{\Psi}}}$ exactly satisfies $\boldsymbol{\mathrm{\Psi}}\boldsymbol{\mathrm{\Psi}}^{\top}=\boldsymbol{\mathrm{I}}$, and the error matrix satisfies $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{\Psi}}}\|\in O(\operatorname{poly}(n)\textbf{{u}})$. Moreover, $\boldsymbol{\mathrm{\Psi}}\begin{pmatrix}\widetilde{\boldsymbol{\mathrm{R}}}\\ 0_{m-n\times n}\end{pmatrix}=\widehat{\boldsymbol{A}}$, where $\widehat{\boldsymbol{A}}=\boldsymbol{\mathrm{A}}^{\prime}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{A}}}$ and $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{A}}}\|\leq O(\operatorname{poly}(n)\textbf{{u}})\|\boldsymbol{\mathrm{A}}^{\prime}\|\leq O(\operatorname{poly}(n)\textbf{{u}})$. Note that $\|\widetilde{\boldsymbol{\mathrm{R}}}\|\leq\|\widehat{\boldsymbol{A}}\|\leq\|\boldsymbol{\mathrm{A}}^{\prime}\|+\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{A}}}\|\leq 1+O(\operatorname{poly}(n)\textbf{{u}})$. The total cost of the algorithm is $O(mn^{\omega-1})$ floating point operations (ignoring the negligible term $\eta>0$ in the exponent).
@@
The second result is for the computation of the spectral norm of the matrix from [MMS18] using the Lanczos algorithm.
Theorem A.7 (Imported variant of Theorem 18 from the full version of [MMS18]).
@@
Let $\boldsymbol{\mathrm{B}}\in\mathbb{C}^{m\times n}$. We can compute a value $\widetilde{\Sigma}\in\Theta(\|\boldsymbol{\mathrm{B}}\|)$ with probability at least $1-2/\delta$ in $O\left(mn\log(n)\log(1/\delta)\right)$ floating point operations using $O(\log(n))$ bits of precision.
@@
Let $\boldsymbol{\mathrm{R}}$ be the real part of $\boldsymbol{\mathrm{B}}$ and $\boldsymbol{\mathrm{Z}}$ be the imaginary part. We can approximate $R\in[\tfrac{9}{10}\|\boldsymbol{\mathrm{R}}\|,\|\boldsymbol{\mathrm{R}}\|]$ and $Z\in[\tfrac{9}{10}\|\boldsymbol{\mathrm{Z}}\|,\|\boldsymbol{\mathrm{Z}}]$ using the algorithm of Theorem A.7. We then set $\widetilde{\Sigma}=R+Z\in[0.9(\|\boldsymbol{\mathrm{R}}\|+\|\boldsymbol{\mathrm{Z}}\|,\|\boldsymbol{\mathrm{R}}\|+\|\boldsymbol{\mathrm{Z}}\|]$. Using the triangle inequality we have that
@@ A.6 Symmetrization
for some constant $c$. The same holds for the lower triangular part $\Delta_{L}(\boldsymbol{\mathrm{A}})$. For the diagonal part $\operatorname{diag}(\boldsymbol{\mathrm{A}}))$ it holds that $\|\operatorname{diag}(\boldsymbol{\mathrm{A}})\|\leq\|\boldsymbol{\mathrm{A}}\|$ (see [Bha00] for proofs). Let $\boldsymbol{\mathrm{C}}^{\prime\prime}=\mathsf{HERM}(\boldsymbol{\mathrm{C}}^{\prime})$. By the definition of $\mathsf{HERM}(\cdot)$, the matrix $\boldsymbol{\mathrm{C}}^{\prime\prime}$ can be written as $\boldsymbol{\mathrm{C}}^{\prime\prime}=\Delta_{U}(\boldsymbol{\mathrm{C}}^{\prime})+\Delta_{U}(\boldsymbol{\mathrm{C}}^{\prime})^{*}-\operatorname{diag}(\boldsymbol{\mathrm{C}}^{\prime})=\boldsymbol{\mathrm{C}}+\left(\Delta_{U}(\boldsymbol{\mathrm{E}})+\Delta_{U}(\boldsymbol{\mathrm{E}})^{*}-\operatorname{diag}(\boldsymbol{\mathrm{E}})\right)$. Therefore
@@ A.7 Existing algorithms for invariant subspaces and the (generalized) eigenproblem
We give a brief overview of some results related to the computation of spectral projectors, eigenvectors, and invariant subspaces for the Hermitian definite generalized eigenproblem. Many more details can be found in standard textbooks and references therein [Dem97, Saa11, GVL13, TE05, Par98], as well as the recent overview of [Sri23] for the complexity of the eigenproblem. For further reading we can highlight some influential works from the enormous bibliography on eigenvalue algorithms and perturbation bounds [DK87, Dem87, BD91, Ste72, Ste73, Ste77, Ste79, Ips97, Nak12, BD98, BMX97, Ips00, JS01, DHT01, AINT17, Pai71, Pai76].
@@
The prominent classic method to solve the Hermitian (generalized or regular) eigenproblem is to use numerically stable Householder transformations to reduce a Hermitian matrix to tridiagonal form, and then apply a tridiagonal shifted QR algorithm to diagonalize the tridiagonal matrix, which gives both the eigenvalues and the eigenvectors. The paramount QR algorithm was originally proposed in [Fra61, Fra62, Kub62], its convergence for the symmetric case was analyzed in [Wil68, DT71] in exact arithmetic, and the non-symmetric was recently analyzed in floating point in [BGVS21, BGVS22a, BGVS22b].
@@
With the aforementioned procedure, the floating point complexity is about $\widetilde{O}(n^{3})$ floating point operations, potentially up to $\operatorname{polylog}(1/\epsilon,1/\operatorname{gap}_{k})$ factors. It is known that for Hermitian definite pencils $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$ one can extend this procedure such that, simultanteously, $\boldsymbol{\mathrm{H}}$ will become (Hermitian) tridiagonal and $\boldsymbol{\mathrm{S}}$ the identity [Cra73]. The Hermitian tridiagonal matrix can be further reduced with similarity transformations to a symmetric tridiagonal matrix, in which case one can apply again tridiagonal symmetric QR. Using standard perturbation bounds as the ones used in this work one can potentially obtain provable forward errors for the eigenvectors, and, ultimately, for the invariant subspaces.
@@
While there exist works that overcome the $\widetilde{O}(n^{3})$ worst-case complexity barrier for some computations related to the eigenproblem, none of them provide end-to-end forward errors for individual invariant subspaces. [PC99] showed that the eigenvalues of a matrix can be computed in $O(n^{\omega})$ arithmetic operations, however, it becomes $O(n^{\omega+1})$ boolean operations in rational arithmetic (up to some other omitted factors), and therefore slower than standard $\widetilde{O}(n^{3})$ floating point eigensolvers. [BOE18] proposed a quantum-inspired method for approximate diagonalization in a backward-error sense, with $O(n^{\omega+1})$ bit complexity. [BGVKS22] proved that backward-approximate diagonalization can be solved in $\widetilde{O}(n^{\omega})$ bit complexity in floating point by using smoothed analysis, improving both the $\widetilde{O}(n^{3})$ classic algorithms for the Hermitian case and the $O(\operatorname{poly}(n/\epsilon))$ algorithm of [ABB+18] for the non-Hermitian case. This was extended for the generalized eigenproblem case in [DDS23]. To obtain forward errors for eigenvectors from the backward-approximate solution, the corresponding bounds that are reported in [BGVKS22, DDS23] require the existence of a minimum eigenvalue gap in the original matrix, while the analysis for invariant subspaces was left as an open problem. Our work completes this analysis, and takes a step even further, by showing that (at least based on existing results), explicit diagonalization is redundant for approximating individual invariant subspaces.
@@
The proof uses standard techniques from holomorphic functional calculus and the properties of the pseudospectrum, but it is stated for completeness. A similar proof, for example, can be found in [BD98] for general matrices $\boldsymbol{\mathrm{A}}$ (i.e. in that proof $\boldsymbol{\mathrm{A}}$ can have complex eigenvalues).

### v4

@@ A.2 Floating point model
We assume a standard floating point model of computation and borrow its axioms from [70, Chapter 2]. There is a fixed number of bits to represent floating point numbers, specifically, one bit is reserved for the sign $s$, $p$ bits are used for the exponent $e$, and $t$ bits are used for the significand $m$.
@@
From [70, Theorem 2.2], for all real numbers $\alpha$ in the normalized range it holds that
@@
Divisions and multiplications with $1$ and $2$ do not introduce errors (for the latter we simply increase/decrease the exponent). We assume that we also have an implementation of $\sqrt{\cdot}$ such that $\boldsymbol{\mathsf{fl}}(\sqrt{\alpha})=(1+\theta)\sqrt{\alpha}$ where $|\theta|\leq\textbf{{u}}$. From [70, Lemma 3.1], we can bound products of errors as
@@
The above can be extended also for complex arithmetic (see [70, Lemma 3.5]), where the bound becomes $|\theta|\leq O(\textbf{{u}})$, but we will ignore the constant prefactor for simplicity.
@@ A.3 Spectral decomposition, pseudospectrum, and eigenvalue bounds
In the sections that follow we need to bound the (forward) errors on the eigenvalues of matrices under perturbations that are introduced due to the finite precision arithmetic. Such bounds can be derived by the classic Bauer-Fike theorem the following (cf. [24, 48] for more details).
There are many bounds in the literature describing the effect of perturbations on the eigenvalues. We summarize some classic results in the following proposition and refer to Bhatia’s monograph for a detailed overview [28].
@@
Hermitian Hermitian $|\lambda_{i}(\boldsymbol{\mathrm{A}})-\lambda_{i}(\boldsymbol{\mathrm{B}})|\leq\|\boldsymbol{\mathrm{A}}-\boldsymbol{\mathrm{B}}\|$ [139]
Hermitian Non-Hermitian $|\lambda_{i}(\boldsymbol{\mathrm{A}})-\lambda_{i}(\boldsymbol{\mathrm{B}})|\leq O(\log(n))\|\boldsymbol{\mathrm{A}}-\boldsymbol{\mathrm{B}}\|$ [81, 117]
@@
Recall some useful properties from the seminal book of Trefethen and Embree [135].
Proposition A.1 (Collective results from [135]).
@@
    $\Lambda_{\epsilon-\|\boldsymbol{\mathrm{E}}\|}(\boldsymbol{\mathrm{M}})\subseteq\Lambda_{\epsilon}(\boldsymbol{\mathrm{M}}+\boldsymbol{\mathrm{E}})\subseteq\Lambda_{\epsilon+\|\boldsymbol{\mathrm{E}}\|}(\boldsymbol{\mathrm{M}})$ ([135, Thm 52.4]),
@@
    Any bounded connected component of $\Lambda_{\epsilon}(\boldsymbol{\mathrm{M}})$ has a nonempty intersection with $\Lambda(\boldsymbol{\mathrm{M}})$, i.e., it contains at least one eigenvalue ([135, Thm 4.3]),
@@
    $\bigcup_{i}D(\lambda_{i},\epsilon)\subseteq\Lambda_{\epsilon}(\boldsymbol{\mathrm{M}})\subseteq\bigcup_{i}D(\lambda_{i},\epsilon\kappa_{\boldsymbol{\mathrm{V}}}(\boldsymbol{\mathrm{M}}))$ ([135, Thms 4.3 and 52.2]).
@@ A.5 Imported floating point algorithms for fast linear algebra
Before we dive into the details, we recall the concept of “fast matrix multiplication”, pioneered by Strassen [134], who showed that two square matrices can be multiplied in $O(n^{\omega})$, where $\omega=\log_{2}7\approx 2.807<3$, in real arithmetic. Since then, the matrix multiplication exponent $\omega$ has been significantly reduced, the record to date being $\omega\leq 2.371552$ [47, 141]. In two seminal works [40, 39], Demmel, Dumitriu, Holtz, and Kleinberg, proved that any fast matrix multiplication algorithm can be executed numerically stably in a floating point machine with almost the same arithmetic complexity as in real arithmetic. They also showed that other problems in Numerical Linear Algebra can be reduced to such matrix multiplications, including inversion, LU and QR factorizations, and solving linear systems of equations. Our algorithms build on the existing results for stable fast matrix multiplication [40], inversion [39], as well as backward-approximate diagonalization and the matrix sign function [15]. We import the corresponding results from the aforementioned works in the following theorems.
Theorem A.2 ($\mathsf{MM}$, stable fast matrix multiplication [40]).
@@
Theorem A.3 ($\mathsf{INV}$, logarithmically-stable fast inversion [39]).
@@
Definition A.3 (Circles of Apollonius, imported from Section 4.1 of [15]).
@@
Theorem A.4 ($\mathsf{SGN}$, imported Theorem 4.9 from [15]).
@@
Theorem A.5 ($\mathsf{DEFLATE}$, imported Theorem 5.3 of [15]).
@@
The algorithm succeeds with probability at least $1-\tfrac{(20n)^{3}\sqrt{\beta}}{\epsilon^{2}}$. The number of arithmetic operations is at most $O(T_{\mathsf{MM}}(n))$, and it requires $O(\log(n)+\log(\tfrac{1}{\epsilon}))$ bits of precision. Internally, the algorithm generates $O(n^{2})$ random numbers to form a $n\times n$ complex Ginibre matrix. [^1] We highlight that our Gaussian sampling oracle in Definition D.1 is slightly different than the one described in [15], but it should be straightforward to adapt the analysis using either of the two definitions.
@@
Lastly, we recall the following two results. The first one is for the QR factorization from [39], which we adapt suitably for our analysis, specifically, to compute an orthonormal basis for the column space of a rectangular matrix.
Theorem A.6 (Basis computation, follows from Section 4 in [39]).
@@
We now use the corresponding algorithm of [39] on $\boldsymbol{\mathrm{A}}^{\prime}$. It returns three matrices: an upper triangular matrix $\widetilde{\boldsymbol{\mathrm{R}}}\in\mathbb{R}^{n\times n}$, a matrix $\boldsymbol{\mathrm{W}}\in\mathbb{R}^{m\times n}$ with $\|\boldsymbol{\mathrm{W}}\|\in O(n)$, and a matrix $\boldsymbol{\mathrm{Y}}\in\mathbb{R}^{n\times m}$ with $\|\boldsymbol{\mathrm{Y}}\|\in O(n)$. The matrix $\boldsymbol{\mathrm{\Psi}}=\boldsymbol{\mathrm{I}}-\boldsymbol{\mathrm{W}}\boldsymbol{\mathrm{Y}}^{\top}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{\Psi}}}$ exactly satisfies $\boldsymbol{\mathrm{\Psi}}\boldsymbol{\mathrm{\Psi}}^{\top}=\boldsymbol{\mathrm{I}}$, and the error matrix satisfies $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{\Psi}}}\|\in O(\operatorname{poly}(n)\textbf{{u}})$. Moreover, $\boldsymbol{\mathrm{\Psi}}\begin{pmatrix}\widetilde{\boldsymbol{\mathrm{R}}}\\ 0_{m-n\times n}\end{pmatrix}=\widehat{\boldsymbol{A}}$, where $\widehat{\boldsymbol{A}}=\boldsymbol{\mathrm{A}}^{\prime}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{A}}}$ and $\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{A}}}\|\leq O(\operatorname{poly}(n)\textbf{{u}})\|\boldsymbol{\mathrm{A}}^{\prime}\|\leq O(\operatorname{poly}(n)\textbf{{u}})$. Note that $\|\widetilde{\boldsymbol{\mathrm{R}}}\|\leq\|\widehat{\boldsymbol{A}}\|\leq\|\boldsymbol{\mathrm{A}}^{\prime}\|+\|\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{A}}}\|\leq 1+O(\operatorname{poly}(n)\textbf{{u}})$. The total cost of the algorithm is $O(mn^{\omega-1})$ floating point operations (ignoring the negligible term $\eta>0$ in the exponent).
@@
The second result is for the computation of the spectral norm of the matrix from [107] using the Lanczos algorithm.
Theorem A.7 (Imported variant of Theorem 18 from the full version of [107]).
@@
Let $\boldsymbol{\mathrm{B}}\in\mathbb{C}^{m\times n}$. We can compute a value $\widetilde{\Sigma}\in\Theta(\|\boldsymbol{\mathrm{B}}\|)$ with probability at least $1-2\delta$ in $O\left(mn\log(n)\log(1/\delta)\right)$ floating point operations using $O(\log(n))$ bits of precision.
@@
Let $\boldsymbol{\mathrm{R}}$ be the real part of $\boldsymbol{\mathrm{B}}$ and $\boldsymbol{\mathrm{Z}}$ be the imaginary part. We can approximate $R\in[\tfrac{9}{10}\|\boldsymbol{\mathrm{R}}\|,\|\boldsymbol{\mathrm{R}}\|]$ and $Z\in[\tfrac{9}{10}\|\boldsymbol{\mathrm{Z}}\|,\|\boldsymbol{\mathrm{Z}}\|]$ using the algorithm of Theorem A.7. We then set $\widetilde{\Sigma}=R+Z\in[0.9(\|\boldsymbol{\mathrm{R}}\|+\|\boldsymbol{\mathrm{Z}}\|),\|\boldsymbol{\mathrm{R}}\|+\|\boldsymbol{\mathrm{Z}}\|]$. Using the triangle inequality we have that
@@ A.6 Symmetrization
for some constant $c$. The same holds for the lower triangular part $\Delta_{L}(\boldsymbol{\mathrm{A}})$. For the diagonal part $\operatorname{diag}(\boldsymbol{\mathrm{A}}))$ it holds that $\|\operatorname{diag}(\boldsymbol{\mathrm{A}})\|\leq\|\boldsymbol{\mathrm{A}}\|$ (see [27] for proofs). Let $\boldsymbol{\mathrm{C}}^{\prime\prime}=\mathsf{HERM}(\boldsymbol{\mathrm{C}}^{\prime})$. By the definition of $\mathsf{HERM}(\cdot)$, the matrix $\boldsymbol{\mathrm{C}}^{\prime\prime}$ can be written as $\boldsymbol{\mathrm{C}}^{\prime\prime}=\Delta_{U}(\boldsymbol{\mathrm{C}}^{\prime})+\Delta_{U}(\boldsymbol{\mathrm{C}}^{\prime})^{*}-\operatorname{diag}(\boldsymbol{\mathrm{C}}^{\prime})=\boldsymbol{\mathrm{C}}+\left(\Delta_{U}(\boldsymbol{\mathrm{E}})+\Delta_{U}(\boldsymbol{\mathrm{E}})^{*}-\operatorname{diag}(\boldsymbol{\mathrm{E}})\right)$. Therefore
@@ A.7 Existing algorithms for invariant subspaces and the (generalized) eigenproblem
We give a brief overview of some results related to the computation of spectral projectors, eigenvectors, and invariant subspaces for the Hermitian definite generalized eigenproblem. Many more details can be found in standard textbooks and references therein [42, 119, 64, 135, 115], as well as the recent overview of [129] for the complexity of the eigenproblem. For further reading we can highlight some influential works from the enormous bibliography on eigenvalue algorithms and perturbation bounds [44, 43, 8, 130, 131, 132, 133, 75, 108, 9, 26, 76, 77, 37, 2, 113, 112].
@@
The prominent classic method to solve the Hermitian (generalized or regular) eigenproblem is to use numerically stable Householder transformations to reduce a Hermitian matrix to tridiagonal form, and then apply a tridiagonal shifted QR algorithm to diagonalize the tridiagonal matrix, which gives both the eigenvalues and the eigenvectors. The paramount QR algorithm was originally proposed in [51, 52, 90], its convergence for the symmetric case was analyzed in [140, 38] in exact arithmetic, and the non-symmetric was recently analyzed in floating point in [18, 16, 17]. Other classical algorithms for computing invariant subspaces include [97, 98, 10, 9].
@@
With the aforementioned procedure, the floating point complexity is about $\widetilde{O}(n^{3})$ floating point operations, potentially up to $\operatorname{polylog}(1/\epsilon,1/\operatorname{gap}_{k})$ factors. It is known that for Hermitian definite pencils $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$ one can extend this procedure such that, simultanteously, $\boldsymbol{\mathrm{H}}$ will become (Hermitian) tridiagonal and $\boldsymbol{\mathrm{S}}$ the identity [36]. The Hermitian tridiagonal matrix can be further reduced with similarity transformations to a symmetric tridiagonal matrix, in which case one can apply again tridiagonal symmetric QR. Using standard perturbation bounds as the ones used in this work one can potentially obtain provable forward errors invariant subspaces.
@@
While there exist works that overcome the $\widetilde{O}(n^{3})$ worst-case complexity barrier for some computations related to the eigenproblem, none of them provide end-to-end forward errors for individual invariant subspaces. [114] showed that the eigenvalues of a matrix can be computed in $O(n^{\omega})$ arithmetic operations, however, it becomes $O(n^{\omega+1})$ boolean operations in rational arithmetic (up to some other omitted factors), and therefore slower than standard $\widetilde{O}(n^{3})$ floating point eigensolvers. [25] proposed a quantum-inspired method for approximate diagonalization in a backward-error sense, with $O(n^{\omega+1})$ bit complexity. [15] proved that backward-approximate diagonalization can be solved in $\widetilde{O}(n^{\omega})$ bit complexity in floating point by using smoothed analysis, improving both the $\widetilde{O}(n^{3})$ classic algorithms for the Hermitian case and the $O(\operatorname{poly}(n/\epsilon))$ algorithm of [7] for the non-Hermitian case. This was extended for the generalized eigenproblem case in [41]. As already mentioned in the introduction, to obtain forward errors for eigenvectors from the backward-approximate solution, the corresponding bounds that are reported in [15, 41] require the existence of a minimum eigenvalue gap in the original matrix, while the analysis for invariant subspaces was left as an open problem. Our work completes this analysis, and takes a step even further, by showing that (at least based on existing results), explicit diagonalization is redundant for approximating individual invariant subspaces.
@@
The proof uses standard techniques from holomorphic functional calculus and the properties of the pseudospectrum, but it is stated for completeness. A similar proof, for example, can be found in [9] for general matrices $\boldsymbol{\mathrm{A}}$ (i.e. in that proof $\boldsymbol{\mathrm{A}}$ can have complex eigenvalues).